OpenStudy (anonymous):

For 0°≤θ≤360°, how many roots does the equation tanx=2sinx have?

OpenStudy (anonymous):

start with writing \(\tan \theta\) as \(\frac{\sin \theta}{\cos \theta}\)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

u droped some of solutions with canceling \(\sin x\)

OpenStudy (anonymous):

sinθ/cosθ = 2sinθ sinθ/2sinθ = cosθ 1/sinθ=cosθ ??

OpenStudy (anonymous):

ok see u have\[\frac{\sin x}{\cos x}=2 \sin x\]multiply by \(\cos x\)\[\sin x=2\sin x \cos x\]\[\sin x-2\sin x \cos x=0\]\[\sin x(1-2 \cos x)=0\]now use zero product property

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

sinx=0 x=0,180,360

OpenStudy (anonymous):

1-2cosx=0 cosx=-1/2 ?

OpenStudy (anonymous):

x=120?

OpenStudy (anonymous):

be careful :) \[\cos x=\frac{1}{2}\]ha? :)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

60?

OpenStudy (anonymous):

just 60 ?

OpenStudy (anonymous):

300?

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

5 roots?

OpenStudy (anonymous):

quite right :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome :)